OpenStudy (anonymous):

Evaluate the following limits. (a) limx→∞(4/(ex−9))=

OpenStudy (anonymous):

please help me!!?

OpenStudy (anonymous):

put x=1/y \[y \rightarrow0\] \[\lim_{x \rightarrow \infty}\frac{ 4 }{ex-9 }=\lim_{y \rightarrow 0}\frac{ 4 }{ \frac{ e }{y }-9 }\] \[=\lim_{y \rightarrow0}\frac{ 4y }{e-9y }=\frac{ 4*0 }{e-9*0 }=\frac{ 0 }{e }=0\]

OpenStudy (anonymous):

what will happen if it goes to the -infinity, if you don't mind me asking @surjithayer

OpenStudy (anonymous):

still the answer remains the same

OpenStudy (anonymous):

If you want to be technical about it, as it goes to infinity we have positive finite / positive infinite => +0. As it goes to negative infinity, we have a positive finite / negative infinity => -0. The answer is still the same, however it is something you might want to consider when evaluating certain limits. For example, if I were to ask the limits of 1/ (4/(e*x−9) or 1/ *your original limit* - without bringing (e*x-9) up (so that 1/4/(e*x-9) becomes (e*x-9)/4) by knowing the previous limit you could pretty much reduce it to 1/0. However, whether that 0 is a +0 or a -0 is relevant here, because the answer to 1/0 will be either +infinite or -infinite accordingly.